OpenStudy (anonymous):

Find the cot of theta if csc theta = sqrt(17)/4 and tan is greater than 0

OpenStudy (anonymous):

\[Find \cot \theta if \csc \theta = \sqrt17/4 and \tan \theta > 0\]

OpenStudy (anonymous):

1/4?

OpenStudy (anonymous):

Can you explain how you got that? My teacher refuses to help.

OpenStudy (anonymous):

first, you need to figure out which quadrant theta is in

OpenStudy (anonymous):

which one is it?

OpenStudy (anonymous):

first?

OpenStudy (anonymous):

yes, what is csc(theta) = r/y, yes? r=? and y =?

OpenStudy (anonymous):

r=1 & y=4? no?

OpenStudy (anonymous):

no, r = sqrt(17) and y = 4. These were given

OpenStudy (anonymous):

oh.. :l duh

OpenStudy (anonymous):

cot(theta) = x/y yes? what's x?

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

how?